Aspire Faculty ID #16538 · Topic: NIMCET 2008 · Just now
NIMCET 2008

If two events $A$ and $B$ such that $P(A')=0.3,\ P(B)=0.5$ and $P(A\cap B)=0.3$, then $P(B\mid A\cup B')$ is

Solution

$P(A)=1-0.3=0.7$ 
 $P(A\cup B')=P(A)+P(B')-P(A\cap B')$ 
$=0.7+0.5-(0.7-0.3)=0.8$ 
 $P(B\cap(A\cup B'))=P(A\cap B)=0.3$ 
 $P(B\mid A\cup B')=\dfrac{0.3}{0.8}=\dfrac38$ 
 Answer: $\boxed{\dfrac38}$

Previous 10 Questions — NIMCET 2008

Nearest first

Next 10 Questions — NIMCET 2008

Ascending by ID
Ask Your Question or Put Your Review.

loading...